# Exploring the interplay between mass-energy equivalence, interactions and entanglement in an optical lattice clock

### v1

@@ Abstract
We propose protocols that probe manifestations of the mass-energy equivalence in an optical lattice clock (OLC) interrogated with spin coherent and entangled quantum states. To tune and uniquely distinguish the mass-energy equivalence effects (gravitational redshift and second order Doppler shift) in such setting, we devise a dressing protocol using an additional nuclear spin state. We then analyze the interplay between photon-mediated interactions and gravitational redshift and show that such interplay can lead to entanglement generation and frequency synchronization. In the regime where all atomic spins synchronize, we show the synchronization time depends on the initial entanglement of the state and can be used as a proxy of its metrological gain compared to a classical state. Our work opens new possibilities for exploring the effects of general relativity on quantum coherence and entanglement in OLC experiments.
@@
Introduction.—Understanding the interplay between quantum mechanics (QM) and general relativity (GR) is a fundamental quest for modern science. Nevertheless up to date, measurements capable of genuinely witnessing this simultaneous interplay have not been realized in tabletop experiments. A push forward towards this milestone is becoming feasible thanks to recent improvements in precision and accuracy of atomic clocks and interferometers, including matter-wave tests of GR [1, 2, 3, 4, 5, 6, 7], the resolution of the gravitational redshift using spatially separated clocks [8, 9], and more recently within a millimeter-scale sample [10, 11]. These developments [12, 1] open up unique opportunities to search for new physics that could help reconcile the seemingly contradictory predictions of QM and GR.
@@
Figure 1: (a) Schematic of an optical lattice clock (OLC) embedded in the curved spacetime (metric $g_{\mu\nu}$) formed by the earth’s gravity. Mass-energy equivalence is the leading order GR correction that translates internal energy difference $\hbar\omega_{0}$ between $|e\rangle$ and $|g\rangle$ states into a difference in the rest mass of an atom $\hbar\omega_{0}/c^{2}$. Such type of correction generates second-order Doppler shift $E_{\rm SDS}$ and gravitational redshift $E_{\mathrm{GRS},j}$ to the clock transition (see Eq. (3) and Eq. (4)). (b) Schematic of the interplay between gravitational redshift and collective cavity-mediated interactions (see Eq. (5)), with $J_{\perp}$ and $J_{z}$ collective exchange and Ising couplings.
@@
Experimental developments have in parallel driven a great deal of theoretical effort towards the understanding of quantum dynamics with GR corrections. These progresses encompass analyses of relativistic corrections to Hamiltonians considered specifically in the context of neutral-atom and trapped-ion systems [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26], tests of mass-energy equivalence with atoms in internal superposition states including predictions of energy-dependent phase shifts, loss of coherence and spin-motion coupling induced by gravitational time dilation [27, 28, 29, 30, 31, 32], among others [33, 34, 35, 36, 37, 38, 39]. However, understanding the direct consequence of GR effects in more complex scenarios such as many-body systems, where particles can interact over the entire array, remains an outstanding problem.
@@
In this work, we provide a first step in this direction by proposing near-term protocols to explore the manifestations of the leading order single-atom GR corrections in quantum many-body dynamics, taking advantage of the state-of-the-art Wannier-Stark OLCs with an array of fully controllable interacting particles under gravity [10, 40]. We perform a detailed analysis of the effects of mass-energy equivalence in a clock transition, i.e. net change in relativistic rest mass due to internal energy, including the gravitational redshift and the second-order Doppler shift. We devise a dressing protocol using an additional nuclear spin state that can be used to tune and uniquely distinguish mass-energy equivalence in an OLC. Taking advantage of the dressing protocol, we further analyze the interplay between photon-mediated interactions and the gravitational redshift acting on individual atoms. While gravitational redshift is akin to a mere level shift due to a magnetic field gradient, the observation in future OLCs will provide a direct test of GR effects acting on a many-body system. Depending on the relative strength of the redshift compared to a many-body energy gap induced by interactions, the system can feature local or global synchronization, as well as entanglement generation for both cases. For the latter case, we show that the synchronization time depends on the entanglement of the initial state and can be used as a proxy for the state’s metrological utility.
@@
Mass-energy equivalence in OLCs.—We consider a single atom in earth’s gravity described by a curved spacetime metric $g_{\mu\nu}$ (see Fig. 1(a)). We perform post-Newtonian expansion of $g_{\mu\nu}$ in power series of $\phi/c^{2}$ [41, 42], with $\phi\approx gZ$ the Newtonian gravitational potential near the earth’s surface and $g$ the local gravitational acceleration. Following the treatment in Ref. [13, 16, 18, 21], one can obtain a single-atom Hamiltonian $\hat{H}_{A}$ accounting for the leading relativistic corrections:
@@
Here, $\hat{H}_{\mathrm{point}}(M)=Mc^{2}+\hat{H}_{0}(M)+\hat{H}_{\mathrm{other}}$ is the Hamiltonian of a point particle with mass $M$, $\hat{H}_{0}(M)=\hat{\mathbf{P}}^{2}/(2M)+M\phi$ contains the non-relativistic terms, and $\hat{H}_{\mathrm{other}}$ contains other GR corrections arising from motion and metric geometry corrections  [^1] See Supplemental Material at [URL will be inserted by publisher] for details of post-Newtonian corrections in an OLC, dressed state protocols, additional numerical simulations of superexchange interactions, as well as analytical derivation of synchronization time, includes Ref. [41, 18, 21, 26, 13, 10, 40, 16, 44, 56, 57, 58, 59, 60, 61, 62]. The key idea of Eq. (1) can be understood as the mass-energy equivalence, summarized by the replacement $M\to M+\hat{H}_{I}/c^{2}$ in $\hat{H}_{\mathrm{point}}$. OLCs feature an ultranarrow optical transition (clock transition) between two long-lived electronic states (excited state $|e\rangle$, ground state $|g\rangle$), which is described by the internal Hamiltonian $\hat{H}_{I}=\hbar\omega_{0}|e\rangle\langle e|$, with $\omega_{0}$ the clock transition frequency measured at the lab position $Z=0$ (see Fig. 1(a)). Since in an OLC $\hat{H}_{I}$ contains the largest observable energy scale compared to other terms, the mass-energy equivalence is the leading order GR correction. It translates into a difference in the rest mass of an atom in states $|e\rangle$ and $|g\rangle$: $M_{g}=M$, $\Delta M=M_{e}-M_{g}=\hbar\omega_{0}/c^{2}$. Note that the mass defect $\Delta M$ is not simply a fixed number, and its tunability (see Fig. 2) is an important tool to determine the relativistic origin of the mass defect.
We assume that in OLCs atoms are trapped in a magic-wavelength 1D lattice along the gravitational potential ($Z$ axis), where $|e\rangle$ and $|g\rangle$ states experience the same potential, $V(Z)=V_{Z}\sin^{2}(k_{L}Z)+M\phi$ [10, 40]. Here $V_{Z}$ is the lattice depth, $k_{L}$ is the wave number of the lattice that sets the atomic recoil energy $E_{R}=\hbar^{2}k_{L}^{2}/2M$ and lattice spacing $a_{L}=\pi/k_{L}$. Metric geometry can also lead to corrections to optical lattices [26], which are negligible in our case. In a tilted 1D lattice described by $V(Z)$, the motional eigenstates in the ground band are the so-called Wannier-Stark (WS) states $|W_{j}\rangle$, with $j$ the $Z$-lattice site index where the WS state is centered at. Assuming the radial degrees of freedom are also confined to the lowest ground state by an additional 2D lattice, with lattice depths $V_{X,Y}$, the eigenenergies of WS states are given by $E_{j}=Mga_{L}j+E_{\mathrm{band}}$, where $E_{\mathrm{band}}\approx\sum_{\alpha=X,Y,Z}E_{R}(\sqrt{V_{\alpha}/E_{R}}-1/4)$ is the ground band zero-point energy. The GR corrections due to mass-energy equivalence is given by
@@
Applying the mass-energy equivalence to the gravitational potential energy $M\phi$, we get the GRS (or gravitational time dilation),
$E_{\mathrm{GRS},j}=\frac{\Delta M}{M}\langle W_{j}|M\phi|W_{j}\rangle=\hbar\omega_{0}\frac{ga_{L}j}{c^{2}}.$ (3)
The GRS leads to a gradient of frequency shift across the lattice. For example, the fractional frequency difference for nearest-neighbor lattice sites is just $4.4\times 10^{-23}$, while it is at the order of $10^{-19}$ for $1$ mm spatial separation as recently observed [10, 11].
The contribution of mass-energy equivalence in the kinetic energy leads to a local modification of the zero-point energy known as SDS (or motional time dilation),
@@
Figure 2: Tuning mass-energy equivalence via dressed states. (a) Schematic of dressing the clock transition with another nuclear spin. The left panel show the application of the dressing laser, and the right panel show the new clock transition in the dressed basis. (b) The tunability of the mass defect $\Delta M$ and gravitational redshift $\omega_{\mathrm{GRS}}$ as a function of dressing parameter $\delta/\Omega$. $\Delta M_{0}$ and $\omega_{\mathrm{GRS},0}$ are the corresponding values without dressing. (c) The tunability of cavity-mediated interactions (see $\hat{H}_{\rm cav}$ in Eq. (5)) as a function of dressing parameters $\delta/\Omega$ and nuclear spin level $m_{F}$. Heisenberg interaction ($J_{\perp}=J_{z}$) can be reached with $m_{F}=3/2$.
Tuning and distinguishing GR effects.—In standard OLCs, the effects of GRS might be mimicked by a weak magnetic field gradient. To observe genuine GR effects, one approach is to simultaneously observe $E_{\mathrm{GRS},j}$ and $E_{\rm SDS}$ in the same system. This should be possible in next-term OLCs by populating higher motional bands if the systematic uncertainty of lattice Stark shifts [43] is suppressed below $10^{-20}$.
@@
This protocol is feasible thanks to the fact that the clock transitions between different nuclear spins are frequency resolved due to magnetic Zeeman shifts. We also assume all other dynamical frequencies are smaller than the dressed state energy gap $\sqrt{\Omega^{2}+\delta^{2}}$ and the Zeeman shifts between nuclear spins. To guarantee the matching of laser phases for each atom, the dressing beam and the clock beam should be co-propagating. Moreover, spatial inhomogeneities in atomic detunings $\delta$ and in the dressing laser Rabi frequency $\Omega$ might obscure the effects of gravitational redshift. For a mHz gravitational redshift arising from a cm-scale spatial separation, it would be required to suppress the spatial variations of $\delta(Z)$, $\Omega(Z)$ and other source of perturbations below $10^{-4}$ Hz. This has already been achieved in Ref. [10] except the requirement for $\Omega(Z)$, which could be achievable using a cavity to stabilize the spatial mode of the dressing laser. One can also circumvent the stringent requirement for $\Omega(Z)$ by averaging the transition frequency of $|g,m_{F}\rangle\leftrightarrow|-\rangle$ and $|g,m_{F}\rangle\leftrightarrow|+\rangle$, while sacrificing the tunability of mass defect ($\Delta M=\Delta M_{0}/2$) [42].
@@
Interplay with many-body dynamics.—After providing a recipe to distinguish genuine GR effects in OLCs, we further explore their manifestations in quantum many-body dynamics. We consider photon-mediated interactions generated by placing the atoms in an optical cavity [44, 45], in a regime where atomic contact interactions are controlled to be much weaker. The interplay between photon-mediated interactions and the GRS is described by the following Hamiltonian (see Fig. 1(b)):
$\hat{H}_{\rm cav}/\hbar=J_{\perp}\hat{\bf S}\cdot\hat{\bf S}+(J_{z}-J_{\perp})\hat{S}^{z}\hat{S}^{z}+\omega_{\mathrm{GRS}}\sum_{j}j\hat{\mathcal{S}}^{z}_{j},$ (5)
@@
where $\hbar\omega_{\mathrm{GRS}}=(\Delta M)ga_{L}$ is the GRS between nearest neighbor sites, $J_{\perp}$ and $J_{z}$ are the collective exchange and Ising couplings. Here $\hat{\mathcal{S}}_{j}^{x,y,z}$ are collective spin operators summed over all atoms at the same height $ja_{L}$, and $\hat{S}^{x,y,z}=\sum_{j}\hat{\mathcal{S}}_{j}^{x,y,z}$. Based on Eq. (5), a magnetic field gradient will in principle give rise to similar single-atom inhomogeneities in the Hamiltonian. We drop the GR corrections for interaction terms since they are negligible in our case [42]. While the use of a single nuclear spin state restricts the cavity exchange interactions to a single polarization mode ($J_{z}=0$ only), the dressing to another nuclear spin allows for coupling two polarization modes of the cavity [42], which enhances the tunability of $\hat{H}_{\rm cav}$ and realizes collective Heisenberg interactions ($J_{\perp}=J_{z}$) as shown in Fig. 2(c). In the following, we mainly focus on the case of $J_{\perp}=J_{z}$, since the $\hat{\bf S}\cdot\hat{\bf S}$ term becomes a constant and does not alter entanglement in the fully symmetric manifold. This requirement is unnecessary for observing frequency synchronization.
@@
We propose to initialize all the atoms in the state $(|g\rangle+|e\rangle)/\sqrt{2}$, perform time evolution under Hamiltonian $\hat{H}_{\rm cav}$ (Eq. (5)), and then measure the phase shift of $\langle\hat{\mathcal{S}}^{+}_{j}\rangle$ for every $Z$-lattice site, resulting in frequency shift $\omega_{j}(t)=\tan^{-1}\left[\langle\hat{\mathcal{S}}^{y}_{j}\rangle/\langle\hat{\mathcal{S}}^{x}_{j}\rangle\right]/t$ as a function of evolution time (see Fig. 3(a)). It can be observed by the application of a $\pi/2$ pulse followed by local population measurements. Without interactions or in the case of short interrogation times, one expects to observe the GRS, $\omega_{j}=j\omega_{\mathrm{GRS}}$, as reported in Ref. [10]. In the interaction dominated regime, the GRS persists only for a time scale shorter than the atomic interaction time scale. Beyond this period, the frequencies become synchronized due to interaction locking and reach $\omega_{j}\approx 0$ at synchronization time $t_{\mathrm{syn}}$ (see Fig. 3(b)). Without loss of generality, the numerical simulations in Fig. 3 and Fig. 4 are based on exact diagonalization for $N=16$ atoms with one atom per $Z$-lattice site.
Figure 3: Interplay between photon-mediated interactions and GRS. (a) We prepare a product state with all atoms in $|g\rangle$ state and apply a laser pulse $\hat{R}_{y}^{-\pi/2}=\exp(i\frac{\pi}{2}\hat{S}^{y})$ to start the dynamics. We focus on a single chain with $N=16$ atoms under the Hamiltonian $\hat{H}_{\rm cav}$ (Eq. (5)) with $J_{\perp}=J_{z}$. (b) Individual atomic frequency shift $\omega_{j}$ with $\omega_{\rm split}/NJ_{\perp}=0.3125$. Synchronization of atomic frequencies can be reach at time $t_{\mathrm{syn}}$. (c) Individual atomic frequency shift $\omega_{j}$ with $\omega_{\rm split}/NJ_{\perp}=3.125$. Global synchronization fails to occur in this regime. (d) Spin squeezing parameter $\xi^{2}$ and normalized Rényi entropy $\tilde{S}_{N/2}$ (inset) in the case of (b). (e) Normalized Rényi entropy $\tilde{S}_{N/2}$ in the case of (c).
@@
Figure 4: Interplay between entanglement and GRS. (a) One-axis twisting (OAT) interactions $\hat{U}_{\mathrm{OAT}}(Q)=\exp(-iQ\hat{S}^{z}\hat{S}^{z}/N)$ and rotations $\hat{R}_{x}^{\theta}=\exp(-i\theta\hat{S}^{x})$ are first applied to generate a spin squeezed initial state (at $Q/2\pi=0.6$), followed by unitary evolution under $\hat{H}_{\rm cav}$. (b) The synchronization time, $t_{\mathrm{syn}}$ depends on the orientation of the spin squeezed state determined by $\theta$. (c) $\Delta t_{\mathrm{syn}}$ (marked in (b)) as a function of OAT shearing strength $Q$. We show $4\mathrm{Cov}(y,z)$ is approaching the quantum Fisher information $F_{Q}$ for spin squeezed states (inset). We compare the numerical simulations (blue points) for $N=16$ atoms under $\hat{H}_{\rm cav}$ ($J_{\perp}=J_{z}$) with Eq. (6) (black dashed lines).
@@
The emergent synchronization is the result of many-body gap protection also observed in prior experiments [46, 47, 44, 48, 49, 50], which arises when $\omega_{\rm split}\ll\Delta E$. Here $\omega_{\rm split}=(N_{s}-1)\omega_{\rm GRS}$ is the maximum redshift in the array, with $N_{s}$ the number of lattice sites, and $\Delta E$ is the many-body gap due to Heisenberg couplings. On the contrary, in the regime $\omega_{\rm split}\sim\Delta E$, the gap cannot maintain global synchronization (see Fig. 3(c)). Using a spin wave analysis one obtains $\Delta E=NJ_{\perp}$ and $NJ_{\perp}t_{\mathrm{syn}}=\pi$ for $J_{\perp}=J_{z}$, where $N$ corresponds to the total atom number in the array. For $N\sim 10^{4}-10^{5}$ 87Sr atoms [44], one can achieve $NJ_{\perp}/2\pi\sim$ Hz ($\omega_{\rm split}/NJ_{\perp}\sim 10^{-3}$ for cm-scale separation), leading to an achievable synchronization time scale in experiment.

### v2

# Exploring the dynamical interplay between mass-energy equivalence, interactions and entanglement in an optical lattice clock

@@ Abstract
We propose protocols that probe manifestations of the mass-energy equivalence in an optical lattice clock (OLC) interrogated with spin coherent and entangled quantum states. To tune and uniquely distinguish the mass-energy equivalence effects (gravitational redshift and second order Doppler shift) in such a setting, we devise a dressing protocol using an additional nuclear spin state. We then analyze the dynamical interplay between photon-mediated interactions and gravitational redshift and show that such interplay can lead to entanglement generation and frequency synchronization dynamics. In the regime where all atomic spins synchronize, we show the synchronization time depends on the initial entanglement of the state and can be used as a proxy of its metrological gain compared to a classical state. Our work opens new possibilities for exploring the effects of general relativity on quantum coherence and entanglement in OLC experiments.
@@
Introduction.—Understanding the interplay between quantum mechanics (QM) and general relativity (GR) is a fundamental quest for modern science. Nevertheless to date, measurements capable of genuinely witnessing this simultaneous interplay have not been realized in tabletop experiments. A push forward towards this milestone is becoming feasible thanks to recent improvements in precision and accuracy of atomic clocks and interferometers. For example, matter-wave interferometers have been used for stringent tests on the equivalence principle [1, 2, 3, 4]. In parallel, the resolution of the gravitational redshift using spatially separated clocks has improved from meter scales [5, 6] to sub millimeters [7, 8]. Furthermore, coherence times for both clock and matter-wave interferometer are now sufficiently long [7, 9] to consider integrating clock and interferometer on a single platform. These developments open up unique opportunities to search for new physics [10, 1] that could help reconcile the seemingly contradictory predictions of QM and GR.
@@
Figure 1: (a) Schematic of an optical lattice clock (OLC) embedded in the curved spacetime (metric $g_{\mu\nu}$) formed by the earth’s gravity. Mass-energy equivalence is the leading order GR correction that translates internal energy difference $\hbar\omega_{0}$ between $|e\rangle$ and $|g\rangle$ states into a difference in the rest mass of an atom $\hbar\omega_{0}/c^{2}$. Such type of correction generates second-order Doppler shift $E_{\rm SDS}$ and gravitational redshift $E_{\mathrm{GRS},j}$ to the clock transition (see Eq. (3) and Eq. (4)). (b) Schematic of probing the dynamical interplay between gravitational redshift and collective cavity-mediated interactions (see Eq. (5)).
@@
Experimental developments have in parallel driven a great deal of theoretical effort towards the understanding of quantum dynamics with GR corrections. These progresses encompass analyses of relativistic corrections to Hamiltonians considered specifically in the context of neutral-atom and trapped-ion systems [11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22], tests of mass-energy equivalence with atoms in internal superposition states including predictions of energy-dependent phase shifts, loss of coherence and spin-motion coupling induced by gravitational time dilation [23, 24, 25, 26, 27, 28, 29, 30, 31], among others [32, 33, 34, 35, 36, 37, 38]. However, understanding the direct consequence of GR effects in more complex scenarios such as many-body systems, where particles can interact over the entire array, remains an outstanding problem.
@@
In this work, we provide a first step in this direction by proposing near-term protocols to explore manifestations of single-atom GR effects, including the gravitational redshift (GRS) and the second-order Doppler shift (SDS), in the quantum many-body dynamics of an atomic ensemble. We take advantage of the state-of-the-art Wannier-Stark OLCs that interrogate large arrays of interacting particles under gravity [7, 39]. We analyze the GRS and the SDS specifically in Wannier-Stark OLCs, both governed by the same mechanism known as mass-energy equivalence. To distinctly characterize these GR effects and overcome the limitation that the gravitational redshift acts just as an effective magnetic field gradient in current OLCs, we firstly devise a dressing protocol using an additional nuclear spin state to tune and uniquely distinguish the mass-energy equivalence. Secondly, we propose a spectroscopic protocol to probe the dynamical modification of the GRS due to photon-mediated interactions in the array. Depending on the relative strength of the GRS to the photon-mediated interactions, the atomic phases dynamically evolve from their individual values, dictated by the local GRS, into a semi-local or global synchronized regime. In both cases we observe entanglement growth from an initial product state due to the interplay of the GRS and interactions, in contrast to the entanglement generated directly by gravitational interactions between massive objects [32, 33, 36, 37, 38]. Lastly, we analyze the interplay of the GRS and interactions when they act on an initially entangled state. This is achieved by studying the dependence of the global synchronization time on the entanglement content of the initial state. Remarkably we find the synchronization time can be used as a probe of the state’s metrological utility.
@@
Mass-energy equivalence in OLCs.—We consider a single atom in earth’s gravity described by a curved spacetime metric $g_{\mu\nu}$ (see Fig. 1(a)). We perform an expansion of $g_{\mu\nu}$ in power series of $\phi/c^{2}$ [40, 41], with $\phi(Z)\approx g_{\rm LO}Z$ the Newtonian gravitational potential near the earth’s surface and $g_{\rm LO}$ the local gravitational acceleration. Following the treatment in Ref. [11, 14, 16, 18], one can obtain a single-atom Hamiltonian $\hat{H}_{A}$ accounting for the leading relativistic corrections:
@@
Here, $\hat{H}_{\mathrm{point}}(M)=Mc^{2}+\hat{H}_{0}(M)+\hat{H}_{\mathrm{other}}$ is the Hamiltonian of a point particle with mass $M$, $\hat{H}_{0}(M)=\hat{\mathbf{P}}^{2}/(2M)+M\phi$ contains the non-relativistic terms, and $\hat{H}_{\mathrm{other}}$ contains other GR corrections in center-of-mass coordinates negligible in our case (see  [^1] See Supplemental Material at [URL will be inserted by publisher] for details of GR corrections in an OLC, additional numerical simulations of superexchange interactions, as well as analytical derivation of synchronization time, including Ref. [40, 16, 18, 22, 14, 56, 57, 58, 59, 60, 44]). The key idea of Eq. (1) can be understood as the mass-energy equivalence, summarized by the replacement $M\to M+\hat{H}_{I}/c^{2}$ in $\hat{H}_{\mathrm{point}}$. OLCs feature an ultranarrow optical transition (clock transition) between two long-lived electronic states (excited state $|e\rangle$, ground state $|g\rangle$), which is described by the internal Hamiltonian $\hat{H}_{I}=\hbar\omega_{0}|e\rangle\langle e|$, with $\omega_{0}$ the clock transition frequency measured at the lab position $Z=0$ (see Fig. 1(a)). Since in an OLC $\hat{H}_{I}$ contains the largest observable energy scale compared to other terms, the mass-energy equivalence is the leading order GR correction. It translates into a difference in the rest mass of an atom in states $|e\rangle$ and $|g\rangle$: $M_{g}=M$, $\Delta M_{0}=M_{e}-M_{g}=\hbar\omega_{0}/c^{2}$. Note that in general the mass defect $\Delta M=\langle\hat{H}_{I}\rangle/c^{2}$ is not simply a fixed number, and its tunability (see Fig. 2) is an important tool to determine the relativistic origin of the mass defect.
We assume that in OLCs atoms are trapped in a magic-wavelength 1D lattice along the gravitational potential ($Z$ axis) [7, 39], where $|e\rangle$ and $|g\rangle$ states experience the same potential, $V(Z)=V_{Z}\sin^{2}(k_{L}Z)+M\phi$. Here $V_{Z}$ is the lattice depth, $k_{L}$ is the wave number of the lattice that sets the atomic recoil energy $E_{R}=\hbar^{2}k_{L}^{2}/2M$ and lattice spacing $a_{L}=\pi/k_{L}$. GR corrections to the optical lattice potential are negligible in our case (see [41]). In a tilted 1D lattice described by $V(Z)$, the motional eigenstates in the ground band are the so-called Wannier-Stark (WS) states $|W_{j}\rangle$, with $j$ the $Z$-lattice site index where the WS state is centered at [39]. Assuming the radial degrees of freedom are also confined to the lowest ground state by an additional 2D lattice, with lattice depths $V_{X,Y}$, the eigenenergies of WS states are given by $E_{j}=Mg_{\rm LO}a_{L}j+E_{\mathrm{band}}$, where $E_{\mathrm{band}}\approx\sum_{\alpha=X,Y,Z}E_{R}(\sqrt{V_{\alpha}/E_{R}}-1/4)$ [42] is the ground band zero-point energy. The GR corrections due to mass-energy equivalence is given by
@@
Applying the mass-energy equivalence to the gravitational potential energy $M\phi$, we get the GRS (or gravitational time dilation) between $|e\rangle$ and $|g\rangle$ states ($\Delta M=\Delta M_{0}$),
$E_{\mathrm{GRS},j}=\frac{\Delta M}{M}\langle W_{j}|M\phi|W_{j}\rangle=\hbar\omega_{0}\frac{g_{\rm LO}a_{L}j}{c^{2}}.$ (3)
The GRS leads to a gradient of frequency shifts across the lattice. For example, the fractional frequency difference for nearest-neighbor lattice sites is just $4.4\times 10^{-23}$, while it is at the order of $10^{-19}$ for $1$ mm spatial separation as recently observed [7, 8].
The contribution of mass-energy equivalence in the kinetic energy leads to a local modification of the zero-point energy known as SDS (or motional time dilation in special relativity) between $|e\rangle$ and $|g\rangle$ states,
@@
Figure 2: Tuning mass-energy equivalence via dressed states. (a) Schematic of dressing the clock transition with another nuclear spin. The left panel show the application of the dressing laser, and the right panel show the new clock transition in the dressed basis. (b) The tunability of the mass defect $\Delta M$ and gravitational redshift $\omega_{\mathrm{GRS}}$ as a function of dressing parameter $\delta/\Omega$. $\Delta M_{0}$ and $\omega_{\mathrm{GRS},0}$ are the corresponding values without dressing. (c) The tunability of cavity-mediated interactions (see $\hat{H}_{\rm cGR}$ in Eq. (5)) as a function of dressing parameters $\delta/\Omega$ and nuclear spin level $m_{F}$. Heisenberg interaction ($J_{\perp}=J_{z}$) can be reached with $m_{F}=3/2$.
Tuning and distinguishing GR effects.—In standard OLCs, the effects of GRS might be mimicked by a weak magnetic field gradient. To provide further evidence of genuine GR effects beyond ruling out all possible systematics, one approach is to simultaneously observe $E_{\mathrm{GRS},j}$ and $E_{\rm SDS}$ in the same system. This could be achieved in next-term OLCs by populating higher motional bands if the systematic uncertainty of lattice Stark shifts [43] is suppressed below $10^{-20}$.
@@
This protocol is feasible thanks to the fact that the clock transitions between different nuclear spins are frequency resolved due to magnetic Zeeman shifts. We also assume all other dynamical frequencies are smaller than the dressed state energy gap $\sqrt{\Omega^{2}+\delta^{2}}$ and the Zeeman shifts between nuclear spins. To guarantee the matching of laser phases for each atom, the dressing beam and the clock beam should be co-propagating. Moreover, spatial inhomogeneities in atomic detunings $\delta$ and in the dressing laser Rabi frequency $\Omega$ might obscure the effects of gravitational redshift. For a mHz gravitational redshift arising from a cm-scale spatial separation, the spatial variations of $\delta(Z)$, $\Omega(Z)$ and other source of perturbations need to be suppressed below $10^{-4}$ Hz. The requirement for $\delta(Z)$ is attainable based on the parameters in Ref. [7]. The requirement for $\Omega(Z)$ could be achievable using a cavity to stabilize the spatial mode of the dressing laser. One can also circumvent the stringent requirement for $\Omega(Z)$ by averaging the transition frequency of $|g,m_{F}\rangle\leftrightarrow|-\rangle$ and $|g,m_{F}\rangle\leftrightarrow|+\rangle$ (see Appendix), while sacrificing the tunability of mass defect ($\Delta M=\Delta M_{0}/2$).
@@
Many-body dynamics.— After providing a recipe to distinguish genuine GR effects in OLCs, we further explore their manifestations in quantum many-body dynamics. We consider photon-mediated interactions generated by placing the atoms in an optical cavity [44, 45], in a regime where atomic contact interactions are controlled to be much weaker. The interplay between photon-mediated interactions and the GRS is described by the following Hamiltonian (see Fig. 1(b) and Appendix):
$\hat{H}_{\rm cGR}/\hbar=J_{\perp}\hat{\bf S}\cdot\hat{\bf S}+(J_{z}-J_{\perp})\hat{S}^{z}\hat{S}^{z}+\omega_{\mathrm{GRS}}\sum_{j}j\hat{\mathcal{S}}^{z}_{j},$ (5)
@@
where $\hbar\omega_{\mathrm{GRS}}=(\Delta M)g_{\rm LO}a_{L}$ is the GRS between nearest neighbor sites, $J_{\perp}$ and $J_{z}$ are the collective exchange and Ising couplings. Here $\hat{\mathcal{S}}_{j}^{x,y,z}$ are collective spin operators summed over all atoms at the same height $ja_{L}$, and $\hat{S}^{x,y,z}=\sum_{j}\hat{\mathcal{S}}_{j}^{x,y,z}$. Based on Eq. (5), a magnetic field gradient will in principle give rise to similar single-atom inhomogeneities in the Hamiltonian, but we can tell them apart using the dressing scheme. We drop the GR corrections for interaction terms since they are negligible in our case (see [41]). While the use of a single nuclear spin state restricts the cavity exchange interactions to a single polarization mode ($J_{z}=0$ only), the dressing to another nuclear spin allows for coupling two polarization modes of the cavity (see Appendix), which enhances the tunability of $\hat{H}_{\rm cGR}$ and realizes collective Heisenberg interactions ($J_{\perp}=J_{z}$) as shown in Fig. 2(c). In the following, we mainly focus on the case of $J_{\perp}=J_{z}$, since the $\hat{\bf S}\cdot\hat{\bf S}$ term becomes a constant and does not alter entanglement in the fully symmetric manifold. This requirement is unnecessary for observing frequency synchronization.
@@
We propose to initialize all the atoms in the state $(|\uparrow\rangle+|\downarrow\rangle)/\sqrt{2}$ ($\pi/2$ pulse between the dressed clock levels), perform time evolution under Hamiltonian $\hat{H}_{\rm cGR}$ (Eq. (5)), and then measure the phase shift of $\langle\hat{\mathcal{S}}^{+}_{j}\rangle$ for every $Z$-lattice site, resulting in frequency shift $\omega_{j}(t)=\tan^{-1}\left[\langle\hat{\mathcal{S}}^{y}_{j}\rangle/\langle\hat{\mathcal{S}}^{x}_{j}\rangle\right]/t$ as a function of evolution time (see Fig. 3(a)). It can be observed by the application of a $\pi/2$ pulse followed by local population measurements. Without interactions or in the case of short interrogation times, one expects to observe the GRS, $\omega_{j}=j\omega_{\mathrm{GRS}}$, as reported in Ref. [7]. In the interaction dominated regime, the GRS persists only for a time scale shorter than the atomic interaction time scale. Beyond this period, the frequencies become synchronized due to interaction locking and reach $\omega_{j}\approx 0$ at synchronization time $t_{\mathrm{syn}}$ (see Fig. 3(b)). Without loss of generality, the numerical simulations in Fig. 3 and Fig. 4 are based on exact diagonalization for $N=16$ atoms with one atom per $Z$-lattice site.
[…]
@@
Figure 4: Interplay between entanglement and GRS. (a) One-axis twisting (OAT) interactions $\hat{U}_{\mathrm{OAT}}(Q)=\exp(-iQ\hat{S}^{z}\hat{S}^{z}/N)$ and rotations $\hat{R}_{x}^{\theta}=\exp(-i\theta\hat{S}^{x})$ are first applied to generate a spin squeezed initial state (at $Q/2\pi=0.6$), followed by unitary evolution under $\hat{H}_{\rm cGR}$. (b) The synchronization time, $t_{\mathrm{syn}}$ depends on the orientation of the spin squeezed state determined by $\theta$. (c) $\Delta t_{\mathrm{syn}}$ (marked in (b)) as a function of OAT shearing strength $Q$. We show $C_{yz}=4\mathrm{max}_{\theta}\mathrm{Cov}(y,z)$ is approaching the quantum Fisher information $F_{Q}$ for spin squeezed states (inset). We compare the numerical simulations (blue points) for $N=16$ atoms under $\hat{H}_{\rm cGR}$ ($J_{\perp}=J_{z}$) with Eq. (6) (black dashed lines).
@@
The emergent synchronization is the result of many-body gap protection also observed in prior experiments [46, 47, 44, 48, 49, 50], which arises when $\omega_{\rm split}\ll\Delta E$. Here $\omega_{\rm split}=(N_{s}-1)\omega_{\rm GRS}$ is the maximum redshift in the array, with $N_{s}$ the number of Z-lattice sites, and $\Delta E$ is the many-body gap due to Heisenberg couplings. On the contrary, in the regime $\omega_{\rm split}\sim\Delta E$, the gap cannot maintain global synchronization (see Fig. 3(c)). Using a spin wave analysis one obtains $\Delta E=NJ_{\perp}$ and $NJ_{\perp}t_{\mathrm{syn}}=\pi$ for $J_{\perp}=J_{z}$, where $N$ corresponds to the total atom number in the array [41]. For $N\sim 10^{5}$ 87Sr atoms, one can achieve $NJ_{\perp}/2\pi\sim$ Hz ($\omega_{\rm split}/NJ_{\perp}\sim 10^{-3}$ for cm-scale separation) [44], leading to a synchronization time scale ($\sim 1$s) within reach of current experiments.